OpenStudy (anonymous):

+MEDAL & FAN+ f(x) = 3x – 4 Compare this to the functions f(x) = 2x

OpenStudy (anonymous):

@Nnesha @Needhelpp101 @NotSoGoodAtMath-_- @HelpfulNoona @pooja195 @proud_yemeniah_ @GrandMoffChris @AlexandervonHumboldt2 @imqwerty @isuckatschool43 @Somy @StudyGurl14 @Study312 @Educated @Learner11 @oogieboogie20

OpenStudy (study312):

Try not to spam tags. I can't help with this question.

OpenStudy (anonymous):

Oh shoot it double it stupid computer, thanks...

Nnesha (nnesha):

both equations are in slope intercept form y=mx+b linear equation where m represent the slope (rise over run) and b=y-intercept ( a point where line intersect the y-axis on a graph)

Nnesha (nnesha):

so what is the slope and y-intercept of both functions ?

OpenStudy (anonymous):

(0,-3) & (0,0)?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

???

Nnesha (nnesha):

what is (0,-3) ??

OpenStudy (anonymous):

it's the y-intercept of f(x) = 3^x – 4

Nnesha (nnesha):

`y=mx+b` where b represent y-intercept so what would be the y-intercept of this equation f(x)=3x-4 ?

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

-3?

Nnesha (nnesha):

y=mx+b the number with x variable represent slope not y-intercept constant term in this equation = y-intercept